# High-Capacity and Low-PAPR BICM-OFDM Systems Using Non-Equiprobable and Non-Uniform Constellation Shaping With Clipping and Filtering

### v1

@@ Abstract
We address a design of high-capacity and low-peak-to-average power ratio (PAPR) orthogonal frequency-division multiplexing (OFDM) systems based on bit-interleaved coded modulation (BICM) utilizing non-equiprobable and non-uniform (NENU) constellations as well as clipping and filtering (CAF). The proposed constellations are generated using a truncated Gaussian distribution, and the merging of constellation points, where the former creates a non-uniform constellation (NUC), and the latter decreases the number of signal points without compromising the achievable bit-wise mutual information (BMI). Since the proposed constellations are uniquely determined by only the two parameters, each associated with NUC and cardinality, the complexity required for the numerical optimization process can be significantly low. We focus on the constellation design based on one dimension, i.e., pulse amplitude modulation (PAM), which facilitates the reduction of demapping complexity for the BICM receiver. The use of CAF at the transmitter can efficiently reduce the PAPR of OFDM signals; however, it introduces clipping noise that may degrade error rate performance, making the application of clipping noise cancellation (CNC) at the receiver essential. Therefore, we optimize the NENU constellations in the presence of CAF and CNC. Simulation results demonstrate that the combination of constellation shaping with CAF and CNC enables BICM-OFDM systems to simultaneously achieve low PAPR and high spectral efficiency over additive white Gaussian noise (AWGN) as well as frequency-selective Rayleigh fading channels. Furthermore, comparative studies confirm that the proposed system significantly outperforms the single-carrier counterpart (i.e., DFT-precoded BICM-OFDM) in terms of PAPR and bit error rate (BER) performance over fading channels.
@@ I Introduction
To achieve significantly high spectral efficiency under limited spectral resources, the use of quadrature amplitude modulation (QAM) with very large constellation sizes, together with capacity-approaching channel codes, is essential. A common practical approach is to combine bit-interleaved coded modulation (BICM) and the standard QAM, where the QAM is formed by two orthogonal Gray-labeled and equidistant (i.e., uniform) pulse amplitude modulation (PAM) symbols, each corresponding to in-phase (I) and quadrature (Q) channels. In this manner, coding and modulation can be performed in one-dimensional space. Furthermore, BICM allows us to decompose each space into parallel binary channels, where the conventional binary channel codes are applicable. This structure significantly simplifies the complexity of demapping and decoding processes at the receiver [1].
@@
The major drawback of BICM, especially with the standard Gray-labeled PAM symbols, is its achievable information rates that are inferior to the Shannon limit. The standard uniform PAM constellation, with each constellation point spaced equally from its neighbors, differs significantly from Gaussian signals that reach the Shannon limit. This discrepancy leads to a noticeable gap, especially when aiming for a higher target information rate. To reduce this gap, constellation shaping, which essentially controls the distribution of the PAM constellations to Gaussian, plays an important role.
@@
Constellation shaping is classified as geometric and probabilistic shaping. Probabilistic shaping adjusts the probability distribution of the uniform PAM, whereas geometric shaping modifies the placement of constellation points [2, 3]. Probabilistic shaping is typically implemented by introducing a distribution matcher, which entails high computational and storage complexity at the transmitter [4]. Geometric shaping is achieved by altering the locations of constellation points, making it compatible with current systems [5, 6]. The design of a geometrically shaped constellation, often referred to as a nonuniform constellation (NUC), represents a significant area of ongoing research, particularly in conjunction with BICM [7, 8, 9].
@@
In view of practical applications to systems with high spectral efficiency, we focus on the design for NUC of PAM (i.e., one-dimensional constellation) with a large constellation size. It is well known that the two-dimensional constellations can achieve higher performance in terms of the gap from the Shannon limit, but their implementation becomes challenging as the target information rate increases. In fact, existing wireless communication systems that aim for very high spectral efficiency employ modulation schemes based on one-dimensional constellations. For example, the Advanced Television Systems Committee’s third-generation standard (ATSC 3.0) adopts one-dimensional NUC (PAM) for higher-order cases such as $1024$ and $4096$ constellation points [10].
@@
The major drawback of constellation shaping is that it increases peak-to-average power ratio (PAPR) as it attempts to reduce average power for a given maximum amplitude. It causes a significant reduction of power efficiency at the transmitter, as linear amplification of a high PAPR signal is challenging without sacrificing power amplifier efficiency [11]. This high PAPR issue is especially dominant with one-dimensional shaping, as the shape of the resulting two-dimensional constellation is rectangular rather than circular. Therefore, wireless systems that prioritize power amplifier efficiency, such as satellite communications, adopt circular constellations known as amplitude and phase shift keying (APSK), rather than rectangular constellations due to their lower PAPR. Nevertheless, other modern wireless communication systems adopt orthogonal frequency-division multiplexing (OFDM) for its higher spectral efficiency and robustness against fading channels, and since OFDM transmits multiple signals simultaneously, its PAPR turns out to be high regardless of the constellations used for each subcarrier, with or without shaping [12]. In such a framework, it is necessary to apply PAPR reduction techniques developed for OFDM signals, and they should have a high compatibility with constellation shaping.
@@ I-A Our Contributions
In this work, we design an OFDM system with both high spectral efficiency and low PAPR based on the combination of constellation shaping and clipping and filtering (CAF) [13]. Our constellation shaping is based on the use of one-dimensional NUC, where the number of constellation points is reduced from the original PAM consisting of signal points that are powers of two for reduced complexity without noticeable loss of performance. As a result, some constellation points are selected with higher probability than the others, leading to a non-equiprobable constellation. We refer to the resulting constellations as non-equiprobable and non-uniform (NENU) constellations.
Since CAF causes performance loss due to the in-band nonlinear distortion called clipping noise, we also consider the application of clipping noise cancellation (CNC) [14, 15, 16] such that the degradation is significantly mitigated at the receiver side. Overall, the system can achieve both low PAPR and high spectral efficiency, even within the practical framework of OFDM.
@@
  • We propose a novel design algorithm for NENU PAM constellations, which combines a geometric design based on a truncated Gaussian distribution with a rate-adaptive constellation merging scheme [^1] In [17], we first presented the concept of merging adjacent signal points for constellations designed according to truncated Gaussian distributions. The algorithm presented in this work is a refined version of [17] in that it can enhance the BMI through the merging process. . Numerical evaluations demonstrate that the proposed scheme achieves performance comparable to the numerically optimized NUC adopted in the ATSC standard.
  • We design BICM-OFDM systems with low PAPR and high spectral efficiency, where low PAPR is achieved by CAF at the transmitter and high spectral efficiency is supported by CNC at the receiver that is compatible with our NENU constellations. We optimize the parameters associated with constellation shaping and PAPR reduction by analyzing bit-wise mutual information (BMI) in the context of CAF/CNC.
  • Through computer simulations, we compare the proposed BICM-OFDM systems with the single-carrier counterpart [18] based on discrete Fourier transform (DFT) precoding that employs the same constellation shaping in terms of the complementary cumulative distribution function (CCDF) of PAPR, demonstrating that the proposed system with proper clipping ratio can achieve lower PAPR than the single-carrier systems in the presence of constellation shaping.
  • Comparison of the resulting bit error rate (BER) performance over additive white Gaussian noise (AWGN) and frequency-selective Rayleigh fading channels shows that the proposed BICM-OFDM system provides noticeable gains in both channels. This is due to the CNC that successfully compensates for the BER degradation caused by CAF with no additional penalty in terms of receiver complexity associated with NENU constellations. It outperforms the single-carrier system with the same constellation over fading channels even with lower PAPR.
@@ I-B Related Work
Several systematic geometric shaping approaches based on Gaussian-like QAM constellations have been proposed in the literature [19, 20, 21] to enhance the achievable information rate while keeping design and implementation complexity modest. Despite their respective design variations, these methods share the common design principle of partitioning a Gaussian distribution into equiprobable sub-regions and assigning a representative constellation point to each sub-region, which was originally suggested in [22]. In our previous work [23], we proposed a systematic approach to designing non-uniform QAM constellations using a truncated Gaussian distribution adjusted by a single parameter. This method flexibly generates the constellation ranging from uniform to Gaussian-like distributions of [22], and we identify the parameter that maximizes the BMI under specified channel signal-to-noise ratios (SNRs). Nevertheless, the designed constellations fail to achieve the shaping gains as high as those of ATSC 3.0 [10], which are optimized based on an exhaustive numerical search. Moreover, the underlying design principles of such optimized NUCs remain largely unexplained, limiting theoretical understanding and systematic reproducibility.
@@
As a low-complexity implementation of probabilistic shaping, a many-to-one mapping approach that does not rely on a distribution matcher can be employed [24, 25, 26]. In this method, redundant bit labels are introduced and they are assigned to a smaller number of constellation points such that the resulting probability mass function (PMF) approximates a Maxwell–Boltzmann (i.e., quantized Gaussian) distribution. Extensions to relatively high-order QAM constellations have recently been investigated in [27, 28, 29]. Since these approaches only require modifications to the bit-to-symbol mapping table, their implementation cost is significantly lower than those based on a distribution matcher. However, they tend to result in lower shaping gains due to their coarse approximation of the target distribution unless a considerably large number of shaping bits are used to adjust the PMF. Recent works [30, 31, 32, 33] have also explored the integration of geometric and probabilistic constellation shaping. Many of these approaches design the constellation geometry and symbol probabilities separately through numerical optimization.
In contrast, our approach offers a unified framework that jointly determines both the positions of constellation points and their probabilities based on a systematic design principle. Furthermore, the application of constellation shaping to OFDM systems with reduced PAPR has not been studied in the literature except for some earlier studies based on trellis shaping [34], which generally requires significant computation for PAPR reduction and causes rate loss for constellation shaping associated with the use of shaping bits.
@@ I-C Organization
The rest of this paper is organized as follows: Section II introduces our NENU constellation construction algorithm, followed by the performance analysis with parameter optimization in Section III. Section IV describes the proposed BICM-OFDM system model employing CAF and CNC, where the achievable information rate in this system model is analyzed. Simulation results are presented in Section V, focusing on the BER performance over practical fading channels and the CCDF of the resulting PAPR. Finally, Section VI concludes this work.
@@ I-D Notation
For a given scalar $A_{k}$, the corresponding bold font $\bf{A}$ represents its row vector form. Sets of real positive numbers, complex numbers, and natural numbers are given by $\mathbb{R}$$_{+}$, $\mathbb{C}$}, and $\mathbb{N}$}, respectively, whereas ${\mathbb{F}}_{2}$ represents the binary field; Calligraphy fonts such as $\mathcal{A}$} represent sets, with their cardinality given by $|{\mathcal{A}}|$; ${\cal N}\left(\mu,\sigma^{2}\right)$ and ${\cal CN}\left(\mu,\sigma^{2}\right)$ represent real-valued Gaussian distribution and circularly symmetric complex Gaussian distribution, respectively, both having mean $\mu$ and variance $\sigma^{2}$. Also, $\mathbb{E}\left[\cdot\right]$ denotes the expectation operation.
@@ II A New Constellation Shaping Algorithm
In this section, we develop a novel systematic design for NENU PAM constellations by incorporating a constellation point merging strategy into the truncated Gaussian-based design. While the capacity of coded modulation (CM) systems increases with the number of distinct constellation points, this is not necessarily the case for BICM, as the optimal constellation cardinality that maximizes the BMI depends on the channel SNR. Therefore, we sequentially merge the adjacent Gaussian sub-regions into larger ones according to the target SNR. In other words, we partition a truncated Gaussian distribution into non-equiprobable sub-regions. We demonstrate that when the two design parameters (i.e., the number of distinct signal points and Gaussian truncation threshold) are jointly optimized, the resulting constellations closely match the numerically optimized NUCs without reducing the number of constellation points adopted in the standard. Unlike the latter, which requires a comprehensive table to specify the constellation points, a pair of parameters uniquely determines our constellation.
The proposed constellation design algorithm consists of two primary steps. We first formulate an initial constellation using a truncated Gaussian distribution and then merge the adjacent points. This algorithm thus facilitates a flexible constellation design in terms of the number of constellation points and their distribution, making it suitable for rate-adaptive designs in BICM systems.

### v2

@@ Abstract
We address a design of high-capacity and low-peak-to-average power ratio (PAPR) orthogonal frequency-division multiplexing (OFDM) systems based on bit-interleaved coded modulation (BICM) utilizing non-equiprobable and non-uniform (NENU) constellations as well as clipping and filtering (CAF). The proposed constellations are generated using a truncated Gaussian distribution and the merging of constellation points, where the former creates a non-uniform constellation (NUC), and the latter adjusts the number of signal points for further improving the total bit-wise mutual information (BMI). Unlike other exhaustive search-based approaches, the proposed constellations are uniquely determined by only two parameters associated with NUC and cardinality. Due to this property of limited degrees of freedom, the complexity required for the numerical optimization process can be significantly low. We focus on the constellation design based on one dimension, i.e., pulse amplitude modulation (PAM), which facilitates the reduction of demapping complexity for the BICM receiver. The use of CAF at the transmitter can efficiently reduce the PAPR of OFDM signals; however, it introduces clipping noise that may degrade error rate performance, making the application of clipping noise cancellation (CNC) at the receiver essential. Therefore, we optimize the NENU constellations in the presence of CAF and CNC. Simulation results demonstrate that the combination of constellation shaping with CAF and CNC enables BICM-OFDM systems to simultaneously achieve low PAPR and high spectral efficiency over additive white Gaussian noise (AWGN) as well as frequency-selective fading channels. Furthermore, comparative studies confirm that the proposed system significantly outperforms the single-carrier counterpart (i.e., DFT-precoded BICM-OFDM) in terms of PAPR and bit error rate (BER) performance over fading channels.
@@ I Introduction
To achieve significantly high spectral efficiency under limited spectral resources, the use of quadrature amplitude modulation (QAM) with very large constellation sizes, together with capacity-approaching channel codes, is essential. A common practical approach is to employ bit-interleaved coded modulation (BICM), which separates modulation and coding by decomposing the transmitted constellation into parallel binary-input channels, to which conventional binary channel codes can readily be applied. When combined with standard QAM, the constellation can be decomposed into two Gray-labeled and equidistant (i.e., uniform) pulse amplitude modulation (PAM) constellations, each corresponding to the in-phase (I) and quadrature (Q) components. This structure significantly simplifies the complexity of demapping and decoding processes at the receiver [szczecinski2015bit].
@@
A key challenge in BICM, especially with standard PAM, is its achievable information rates that are inferior to the Shannon limit. The standard PAM constellation, with each constellation point spaced equally from its neighbors, differs significantly from Gaussian signals that achieve the Shannon limit. This discrepancy leads to a noticeable gap, especially when aiming for a high target spectral efficiency. To reduce this gap, constellation shaping, which essentially tailors the distribution of PAM constellations to approximate a Gaussian distribution, plays an important role.
@@
Constellation shaping is categorized as geometric and probabilistic shaping. Probabilistic shaping adjusts the probability distribution of the uniform PAM, whereas geometric shaping modifies the placement of constellation points [steiner2017comparison, qu2019probabilistic]. Probabilistic shaping is typically implemented by introducing a distribution matcher, which entails high computational and storage complexity at the transmitter [schulte2015constant]. In contrast, geometric shaping is simply achieved by altering the locations of constellation points, making it compatible with current systems [fay2016overview, singya2021survey]. Accordingly, the design of a geometrically shaped constellation, often referred to as a non-uniform constellation (NUC), represents a significant area of ongoing research, particularly in conjunction with BICM [sillekens2022high, gumucs2023simplified, chen2022geometrically].
@@
In view of practical applications to systems with high spectral efficiency, we focus on the design of NUCs based on PAM (i.e., one-dimensional NUCs) with large constellation sizes. It is well known that two-dimensional constellations can achieve better performance in terms of the gap from the Shannon limit. However, their implementation becomes increasingly challenging as the target information rate increases due to the prohibitively high demapping complexity associated with large constellation sizes. In fact, existing wireless communication systems targeting very high spectral efficiency often employ modulation schemes based on one-dimensional constellations. For example, the Advanced Television Systems Committee’s third-generation standard (ATSC 3.0) adopts only one-dimensional NUCs for higher modulation orders such as $1024$ and $4096$ [atsc30].
@@
A major drawback of constellation shaping is the increase in peak-to-average power ratio (PAPR), as it aims to reduce the average power for a given maximum amplitude. This leads to a significant reduction of power efficiency at the transmitter, as linear amplification of high PAPR signals is challenging without sacrificing power amplifier efficiency [ochiai2013analysis]. Nevertheless, the majority of modern wireless communication systems adopt orthogonal frequency-division multiplexing (OFDM) because of its advantages, such as high spectral efficiency and robustness against fading channels. Since OFDM transmits multiple signals simultaneously, its PAPR becomes high irrespective of the constellations used for each subcarrier, with or without shaping [ochiai_TCOM_2001]. In such systems, it is necessary to apply PAPR reduction techniques originally developed for OFDM while ensuring high compatibility with constellation shaping.
@@ I-A Our Contributions
In this work, our target is to implement a BICM-OFDM system that achieves both high spectral efficiency and low PAPR without imposing high complexity at the transmitter side.
In our shaping approach, a Gray-labeled non-uniform PAM constellation, whose cardinality is a power of two, is first systematically constructed based on a truncated Gaussian distribution characterized by a single parameter. The number of constellation points is then sequentially reduced by merging the closest points. This merging process improves the achievable information rate at a given signal-to-noise ratio (SNR), while reducing the demapping complexity. As a result, some constellation points are selected with higher probability than the others, leading to a non-equiprobable constellation. We refer to the resulting constellations as non-equiprobable and non-uniform (NENU) constellations.
For achieving low PAPR, we adopt clipping and filtering (CAF) [ochiai2002performance] among various PAPR reduction techniques proposed for OFDM. While many PAPR reduction schemes impose demanding complexity at the transmitter side, CAF can be implemented in a straightforward manner. Furthermore, it is directly applicable to BICM-OFDM systems with any constellation shaping. Therefore, combining constellation shaping with CAF is a natural consequence for BICM-OFDM. However, CAF inevitably causes performance degradation due to the in-band nonlinear distortion called clipping noise. Therefore, we also consider the application of clipping noise cancellation (CNC) [chen2003iterative, sun2021, wachowiak2023clipping] to mitigate this impairment at the receiver. To tailor the constellation design to OFDM transmission under CAF and CNC, the two shaping parameters (i.e., the truncation level associated with the initial Gaussian distribution and the number of distinct signal points) are jointly optimized by taking into account the effects of both additive white Gaussian noise (AWGN) and residual clipping noise. Overall, the proposed system can simultaneously achieve low PAPR and high spectral efficiency, even within a practical framework of OFDM over frequency-selective fading channels.
@@
  • We propose a novel design algorithm for NENU PAM constellations, which combines a geometric design based on a truncated Gaussian distribution with a rate-adaptive constellation merging scheme [^1] In [kurihara2025systematic], we first presented the concept of merging adjacent signal points for constellations designed according to truncated Gaussian distributions. The algorithm presented in this work is a refined version of [kurihara2025systematic] in that it can enhance the BMI through the merging process. . Numerical evaluations demonstrate that the proposed scheme achieves performance comparable to the numerically optimized NUC adopted in the ATSC standard.
  • We design BICM-OFDM systems with low PAPR and high spectral efficiency, where low PAPR is achieved by CAF at the transmitter and high spectral efficiency is supported by CNC at the receiver that is compatible with our proposed shaping scheme. Furthermore, the NENU constellations are optimized in the presence of CAF and CNC.
  • Through computer simulations, we compare the proposed BICM-OFDM systems with their single-carrier counterparts [myung2006single] based on discrete Fourier transform (DFT) precoding that employs the same constellation shaping in terms of the complementary cumulative distribution function (CCDF) of PAPR, demonstrating that the proposed system with the proper clipping ratio can achieve lower PAPR than the single-carrier systems.
  • Comparison of the resulting bit error rate (BER) performance over AWGN and frequency-selective fading channels shows that the proposed BICM-OFDM system provides noticeable gains in both channels. This is due to the CNC that successfully compensates for the BER degradation caused by CAF with no additional complexity associated with the use of constellation shaping. It outperforms the single-carrier system with the same constellation over fading channels even with lower PAPR.
@@ I-B Related Work
Several systematic geometric shaping approaches based on Gaussian-like PAM constellations have been proposed in the literature [koike2017bit, wiens2020constellation, boutros2018geometric] to enhance the achievable information rate while keeping design and implementation complexity modest. Despite their respective design variations, these methods share the common design principle of partitioning a Gaussian distribution into equiprobable subregions and assigning a representative constellation point to each subregion, as originally suggested in [sun1993approaching]. In our previous work [kurihara2024design], we proposed a systematic approach to designing non-uniform PAM constellations using a truncated Gaussian distribution adjusted by a single parameter. This method generates the PAM constellation, whose distribution can be adjusted ranging from Gaussian-like to uniform, and we identify the parameter that maximizes the achievable information rate under specified channel SNRs. Nevertheless, the designed constellations fail to achieve shaping gains as high as those of ATSC 3.0 [atsc30], which rely on the optimization through an exhaustive numerical search. Moreover, the underlying design principles of such optimized NUCs remain largely unexplained, limiting theoretical understanding and systematic reproducibility.
@@
As a low-complexity implementation of probabilistic shaping, a many-to-one mapping approach that does not rely on a distribution matcher can be employed [yankov2014rate, chang2013symbol, zhou2020enhancement]. This approach introduces redundant bit labels, which are assigned to a smaller number of constellation points such that the resulting probability mass function (PMF) approximates a Maxwell–Boltzmann (i.e., quantized Gaussian) distribution. Extensions to relatively high-order QAM constellations have recently been investigated in [yankov2016constellation, jia2023many, cao2020probabilistic]. Since these approaches only require modifications to the bit-to-symbol mapping table, their implementation cost is significantly lower than those based on a distribution matcher. However, they tend to result in lower shaping gains due to their coarse approximation of the target distribution unless a considerably large number of shaping bits are used to adjust the PMF. Recent works [li2024model, soleimanzade2023hybrid, stark2019joint, maneekut2020hybrid] have also explored the integration of geometric and probabilistic constellation shaping. Many of these approaches design the constellation geometry and symbol probabilities separately through numerical optimization.
In contrast, our approach offers a unified framework that jointly determines both the positions of constellation points and their probabilities based on a systematic design principle. We also address the optimization of constellations in the presence of CAF and CNC, which involve complex nonlinear signal processing operations at both the transmitter and receiver. Most existing constellation shaping approaches through numerical optimization deal with analytically tractable channels such as AWGN. On the other hand, since the proposed approach relies on a predefined family of constellations characterized by limited degrees of freedom, it facilitates efficient optimization even for such channels that cannot be described by explicit mathematical models. Lastly, the application of geometric shaping to OFDM systems in conjunction with PAPR reduction techniques has not been well investigated in the literature. [^2] Trellis shaping [forney92] can be applied to PAPR and average power reduction of BICM-OFDM systems [ochiai_comm04, yoshizawa2015]. However, the transmitter complexity required for effective PAPR reduction is considerably higher than CAF, especially when OFDM has a large number of subcarriers.
@@ I-C Organization
The rest of this paper is organized as follows: Section II introduces our NENU constellation construction algorithm, followed by the performance analysis with parameter optimization in Section III. Section IV describes the proposed BICM-OFDM system model employing CAF and CNC, where the achievable information rate in this system model is analyzed. Simulation results are presented in Section V, focusing on the BER performance over AWGN and practical fading channels as well as the CCDF of the resulting PAPR. Finally, Section VI concludes this work.
@@ I-D Notation
For a given scalar $A_{k}$, the corresponding bold font $\mathbf{A}$ denotes its row-vector form. Sets of real positive numbers, complex numbers, and natural numbers are given by $\mathbb{R}$$_{+}$, $\mathbb{C}$}, and $\mathbb{N}$}, respectively, whereas ${\mathbb{F}}_{2}$ represents the binary field. Calligraphic fonts such as $\mathcal{A}$} denote sets, with their cardinality given by $|{\mathcal{A}}|$, whereas ${\cal N}\left(\mu,\sigma^{2}\right)$ and ${\cal CN}\left(\mu,\sigma^{2}\right)$ represent the real-valued Gaussian distribution and the circularly symmetric complex Gaussian distribution, respectively, both having mean $\mu$ and variance $\sigma^{2}$. The expectation operator is represented by $\mathbb{E}\left[\cdot\right]$.
@@ II A New Constellation Shaping Algorithm
The proposed design algorithm for NENU constellations consists of two primary steps. First, an initial Gray-labeled PAM constellation is derived by partitioning a truncated Gaussian distribution into equiprobable subregions, where each constellation point is assigned according to the midpoint cumulative distribution function (CDF) value of the corresponding region. The resulting PAM distribution is controlled by a single parameter that determines the truncation level of the reference Gaussian distribution. Next, adjacent subregions are sequentially merged to adjust the number of distinct signal points. In that case, the reference Gaussian distribution is partitioned such that the probability associated with a given subregion is proportional to the number of assigned binary labels. This algorithm facilitates a flexible constellation design in terms of both the number of constellation points and their distribution, making it suitable for rate-adaptive BICM systems. We demonstrate that when the two design parameters (i.e., the number of distinct signal points and the Gaussian truncation level) are jointly optimized, the resulting constellations closely match the optimized NUCs adopted in the ATSC standard. Unlike the latter, which requires a comprehensive table to specify the constellation points, only a pair of parameters uniquely determines our constellation. As described in Section IV, this “limited degrees of freedom” description turns out to be particularly useful when designing constellations over highly complex channels that may only be modeled by simulation-based approaches.